OpenStudy (anonymous):

5y − 7 = 8 + 4(y − 3)

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

first you distributive property 4(y-3) and you get 4y-12. And the equation now its 5y-7=8+4y-12 then you just combine like terms

OpenStudy (texaschic101):

5y - 7 = 8 + 4(y - 3) 5y - 7 = 8 + 4y - 12 5y - 7 = 4y - 4 5y - 4y = 7 - 4 y = 3 joshhfers is correct :)